et Xi and X2 be independent random variables with common distribution NORM(8,1), where θ is a real number. Let Y = X1 + X a) Find the p.d.f. of Y. 4. L 2. b) Use the appropriate table to find P(Y> 20 + 2)
Answer #1

4)

a)

We know that if X\sim N(\mu,\sigma^{2}) , then its PDF is given by

f(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{\frac{-1}{2 \sigma^{2}}(x-\mu)^{2}},-\infty<X<\infty

Now, X_{i} \sim N(\theta,1)

So, Y=X_{1}+X_{2}\sim N(\theta+\theta,1+1)\sim N(2\theta,2)

Thus PDF of Y is:

g(y)=\frac{1}{\sqrt{2\pi}\sqrt{2}}e^{\frac{-1}{2 (2)}(x-2\theta)^{2}},-\infty<Y<\infty

g(y)=\frac{1}{\sqrt{2\pi}\sqrt{2}}e^{\frac{-1}{4}(x-2\theta)^{2}},-\infty<Y<\infty

b)

Required probability = P(Y>2\theta +2)=P(Z>\frac{(2\theta +2)-2\theta}{\sqrt{2}})=P(Z>1.41)=0.07927